Circular No. 6633

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE P. Garnavich, Harvard-Smithsonian Center for Astrophysics, reports

on behalf of The High-Z Supernova Search Team (IAUC 6160) the discovery of four supernovae on CCD images taken at the Victor Blanco 4-m telescope at Cerro Tololo Interamerican Observatory (CTIO) by N. Suntzeff, R. Schommer, A. Clocchiatti, and M. Phillips (CTIO); and A. Diercks (University of Washington). The supernovae were identified by comparing the images taken in 1997 April with those obtained in 1996.

SN  1997 UT  α_{2000}  δ_{2000}  R z type
1997bu Apr. 3 10^{h}06^{m}09.^{s}99 − 7^{o}16′30′′.1 23 ? ?
1997bv Apr. 3 9 33 48.79 + 0 04 59.4 24 0.43 ?
1997bw Apr. 3 11 30 42.02 − 0 37 56.6 24 0.5 Ia
1997bx Apr. 3 12 25 11.80 + 0 04 54.4 23 0.35 Ia

SNe 1997bu and 1997bv are within 1′′ of their host galaxies. SN 1997bx is 0′′.6 west and 4′′.6 south of its host, while SN 1997bw has no detectable host galaxy. Spectra were obtained with the European Southern Observatory 3.6-m telescope by B. Leibundgut on Apr. 8–9, the Multiple Mirror Telescope by P. Challis on Apr. 7–8, the CTIO 4-m telescope by M. Keane (reduced by R. Covarrubiason) on Apr. 5, and the Keck telescope by A. V. Filippenko, A. G. Riess, and A. J. Barth on Apr. 11. Spectra of some supernovae discovered in March (IAUC 6602) were also observed at the Keck telescope, showing that SN 1997bb is a type-Ia object at z = 0.52 and SN 1997bd is of type Ia at z = 0.67, and confirming that SN 1997as is of type Ia at z = 0.51.

BL CIRCINI Regarding the discussion about BL Cir and GRB 970402 (IAUC 6628,

6632), G. V. Williams, Smithsonian Astrophysical Observatory, writes: “A comparison of the original finder chart (Hoffmeister 1965, A.N. 289, 1) and the Digital Sky Survey (DSS) shows that BL Cir is identical to GSC 9258.2458, an object classified in the Guide Star Catalogue as ‘non-stellar’. The GSC position is therefore suspect. A measurement of the position of BL Cir from the DSS, utilizing 12 USNO A1.0 comparison stars, yields the following J2000.0 position: α = 14^{h}49^{m}56^{s}.68, δ = −69^{o}20′51′′.2 (± 0′′.3), confirming the general correctness of the GCVS position.”

1997 April 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green